# Differentially Private Synthetic Data via APIs 3: Using Simulators Instead of Foundation Models

### v1

@@ 3.3 Sim-PE with Simulator-generated Data
Here, we assume that a dataset of $m$ samples $S_{\text{sim}}=\left\{z_{1},\ldots,z_{m}\right\}$ generated from the simulator is already given. The goal is to pick $N_{\mathrm{syn}}$ samples from them to construct the DP synthetic dataset $S_{\mathrm{syn}}$. Before discussing our final solutions, we first discuss why two straightforward approaches do not work well.
@@
However, the size of the simulator-generated dataset (i.e., $m$) is typically very large (e.g., 1.2 million in Bae et al. (2023)), and the total amount of added Gaussian noise grows with $m$. This means that the resulting histogram suffers from a low signal-to-noise ratio, leading to poor fidelity in $S_{\mathrm{syn}}$.
Baseline 2: Applying DP_NN_HISTOGRAM on cluster centers of $S_{\mathrm{syn}}$. To improve the signal-to-noise ratio of the histogram, one solution is to have private samples vote on the cluster centers of $S_{\text{sim}}$ instead of the raw samples. Specifically, we first cluster the samples in $S_{\text{sim}}$ into $N_{\text{cluster}}$ clusters with centers $\left\{w_{1},\ldots,w_{N_{\text{cluster}}}\right\}$ and have private samples vote on these centers rather than individual samples in $S_{\text{sim}}$. [^1] Note that voting in Lin et al. (2023) is conducted in the image embedding space. Here, $w_{i}$s represent cluster centers in the embedding space, and each private sample uses its image embedding to find the nearest cluster center. Since the number of bins in the histogram decreases from $m$ to $N_{\text{cluster}}$, the signal-to-noise ratio improves. Following the approach of the previous baseline, we then draw $N_{\mathrm{syn}}$ cluster centers (with replacement) based on the histogram and randomly select a sample from each chosen cluster to construct the final $S_{\mathrm{syn}}$.
@@
Our approach. Our key insight is that the unavoidable trade-off between the accuracy of the DP histogram and the precision of selection (clusters vs. individual samples) arises because private samples are forced to consider all samples in $S_{\text{sim}}$—either directly in baseline 1 or indirectly through cluster centers in baseline 2. However, this is not necessary. If we already know that a sample $z_{i}$ is far from the private dataset, then its nearest neighbors in $S_{\text{sim}}$ are also likely to be far from the private dataset. Therefore, we can avoid wasting the privacy budget on evaluating such samples.
@@
Since we only draw $N_{\mathrm{syn}}$ samples (instead of $m$) from RANDOM_API, the DP histogram on this subset has a higher signal-to-noise ratio. In the following steps (Algs. 1, 1 and 1 in Alg. 1), samples far from the private dataset are removed, and we perform variations only on the remaining samples according to:
@@
thus avoiding consideration of nearest neighbors of the removed samples (unless they are also nearest neighbors of retained samples). Similar to $\alpha$ and $\beta$ and in § 3.2, the parameter $\gamma$ controls the degree of variation. At one extreme, when $\gamma=m$, VARIATION_API disregards the input sample and reduces to RANDOM_API. At the other extreme, when $\gamma=1$, VARIATION_API returns the input sample unchanged.
Broader applications. The proposed algorithm can be applied to any public dataset beyond simulator-generated data. In our experiments (§ 4), we focus on simulator-generated data, and we leave the exploration of broader applications for future work.
@@ 3.4 Sim-PE with both Simulators and Foundation Models
As discussed in § 2.1, simulators and foundation models complement each other across different data domains. Moreover, even within a single domain, they excel in different aspects. For example, computer graphics-based face image generation frameworks (Bae et al., 2023; Wood et al., 2021) allow controlled diversity in race, lighting, and makeup while mitigating potential biases in foundation models. However, the generated faces may appear less realistic than those produced by state-of-the-art foundation models. Thus, combining the strengths of both methods for DP data synthesis is highly appealing.
Fortunately, PE naturally supports this integration, as it only requires RANDOM_API and VARIATION_API, which work the same for both foundation models and simulators. While there are many ways to combine them, we explore a simple strategy: using simulators in the early PE iterations to generate diverse seed samples, then switching to foundation models in later iterations to refine details and enhance realism. As shown in § 4, this approach outperforms using either simulators or foundation models alone.
@@ 4.1.1 Datasets and Simulators
Datasets. Following prior work (Gong et al., 2025), we use two private datasets: (1) MNIST (LeCun, 1998), where the image class labels are digits ‘0’-‘9’, and (2) CelebA (Liu et al., 2015), where the image class labels are male and female. We aim at conditional generation for these datasets (i.e., each generated image is associated with the class label).
Simulators. To demonstrate the general applicability of Sim-PE, we select three diverse simulators with very different implementations.
@@
(1) Text rendering program. Generating images with readable text using foundation models is a known challenge (Betker et al., 2023). Simulators can address this gap, as generating images with text through computer programs is straightforward. To illustrate this, we implement our own text rendering program, treating MNIST as the private dataset. Specifically, we use the Python PIL library to render digits as images. The categorical parameters include: (1) Font. We use Google Fonts (Google, 2022), which offers 3589 fonts in total. (2) Text. The text consists of digits ‘0’ - ‘9’. The numerical parameters include: (1) Font size, ranging from 10 to 29. (2) Stroke width, ranging from 0 to 2. (3) Digit rotation degree, ranging from $-30\degree$ to $30\degree$. We set the feasible sets of these parameters to be large enough so that the random samples differ significantly from MNIST (see Fig. 1).
@@
(2) Computer graphics-based renderer for face images. Computer graphics-based rendering is widely used in real-world applications such as game development, cartoons, and movie production. This experiment aims to assess whether these advanced techniques can be adapted for DP synthetic image generation via Sim-PE. We use CelebA as the private dataset and a Blender-based face image renderer from Bae et al. (2023) as the API. Since the source code for their renderer is not publicly available, we apply our data-based algorithm from § 3.3 on their released dataset of 1.2 million face images. It is important to note that this renderer may not necessarily represent the state-of-the-art. As visualized in Fig. 3, the generated faces exhibit various unnatural artifacts and appear less realistic than images produced by state-of-the-art generative models (e.g., Rombach et al. (2022)). Therefore, this experiment serves as a preliminary study, and the results could potentially improve with more advanced rendering techniques.
@@
(3) Rule-based avatar generator. We further investigate whether Sim-PE remains effective when the simulator’s data significantly differs from the private dataset. We use CelebA as the private dataset and a rule-based avatar generator (Escartín, 2021) as the API. This simulator has 16 categorical parameters that control attributes of the avatar including eyes, noses, background colors, skin colors, etc. As visualized in Fig. 4, the generated avatars have a cartoon-like appearance and lack fine-grained details. This contrasts sharply with CelebA images, which consist of real human face photographs.
Class label information from the simulators. For simulator 1, the target class label (i.e., the digit) is fully controlled by one parameter. For simulators 2 and 3, the target class label (i.e., the gender) is not directly controlled by any parameter, but could potentially be obtained by an external image gender classifier. One benefit of using domain-specific simulators is that we can potentially use the class label information to enhance data quality. To get a more comprehensive understanding of Sim-PE, we consider two settings: (1) Class label information is unavailable (abbreviated as “ClassUnavail”). We artificially make the problem more challenging by assuming that the class label information is not available. Therefore, Sim-PE has to learn to synthesize images with the correct class by itself. (2) Class label information is available (abbreviated as “ClassAvail”). On MNIST, we further test how Sim-PE can be improved if the class label information is available. In this case, the RANDOM_API and VARIATION_API (Eqs. 1 and 2) are restricted to draw parameters from the corresponding class (i.e., the digit is set to the target class).
@@ 4.1.2 Metrics and Evaluation Pipelines
We follow the evaluation settings of DPImageBench (Gong et al., 2025), a recent benchmark for DP image synthesis. Specifically, we use two metrics: (1) FID (Heusel et al., 2017) as a quality metric and (2) the accuracy of downstream classifiers as a utility metric. Specifically, we use the conditional version of PE (App. A), so that each generated images are associated with the class labels (i.e., ‘0’-‘9’ digits in MNIST, male vs. female in CelebA). These class labels are the targets for training the classifiers. We employ a strict train-validation-test split and account for the privacy cost of classifier hyperparameter selection. Specifically, we divide the private dataset into disjoint training and validation sets. We then run Sim-PE on the training set to generate synthetic data. Next, we train three classifiers—ResNet (He et al., 2016), WideResNet (Zagoruyko & Komodakis, 2016), and ResNeXt (Xie et al., 2017)—on the synthetic data and evaluate their accuracy on the validation set. Since the validation set is part of the private data, we use the Report Noisy Max algorithm (Dwork et al., 2014) to select the best classifier checkpoint across all epochs of all three classifiers. Finally, we report the accuracy of this classifier on the test set. This procedure ensures that the reported accuracy is not inflated due to train-test overlap or DP violations in classifier hyperparameter tuning.
Following Gong et al. (2025), we set DP parameter $\delta=1/(N_{\mathrm{priv}}\cdot\log N_{\mathrm{priv}})$, where $N_{\mathrm{priv}}$ is the number of samples in the private dataset, and $\epsilon=1$ or 10.
@@ 4.1.3 Baselines
We compare Sim-PE with 12 state-of-the-art DP image synthesizers reported in Gong et al. (2025), including DP-MERF (Harder et al., 2021a), DP-NTK (Yang et al., 2023), DP-Kernel (Jiang et al., 2023), GS-WGAN (Chen et al., 2020), DP-GAN (Xie et al., 2018), DPDM (Dockhorn et al., 2023), PDP-Diffusion (Ghalebikesabi et al., 2023b), DP-LDM (Liu et al., 2024b), DP-LoRA (Tsai et al., 2024), PrivImage (Li et al., 2024), and PE with foundation models (Lin et al., 2023). Except for PE, all other baselines require model training. For Sim-PE with simulator-generated data, we additionally compare it against the two baselines introduced in § 3.3.
It is important to note that this comparison is not intended to be entirely fair, as different methods leverage different prior knowledge. For example, many of the baselines rely on pre-trained foundation models or public datasets from similar distributions, whereas Sim-PE does not. Conversely, Sim-PE utilizes simulators, which none of the baseline methods incorporate. Since Sim-PE is the only approach that leverages simulators, it is more appropriate to consider it as a new evaluation setting or benchmark. The results of other methods serve as a reference point to contextualize this new paradigm among state-of-the-art approaches and to highlight directions for future research.
@@ 4.2 Sim-PE with Simulator Access
In this section, we evaluate Sim-PE with a text rendering program on MNIST dataset. The results are shown in Tables 1, 1, 2 and 2. The key takeaway messages are:
@@
Table 1: Accuracy and FID. The best between PE methods is in bold, and the best between all methods is underlined. “Simulator” refers to samples from the simulator’s RANDOM_API. Results other than Sim-PE and Simulator are taken from Gong et al. (2025).
@@
Quantitative results in Table 1 further support this. Without private data guidance, the simulator naturally generates digits from incorrect classes, leading to a downstream classifier accuracy of only 11.6%, close to random guessing. In contrast, Sim-PE significantly improves accuracy to approximately 90%. Additionally, FID scores confirm that the images generated by Sim-PE more closely resemble real data.
Sim-PE can significantly improve the performance of PE. The PE baseline (Lin et al., 2023) uses a diffusion model pre-trained on ImageNet, which primarily contains natural object images (e.g., plants, animals, cars). Since MNIST differs significantly from such data, PE, as a training-free method, struggles to generate meaningful MNIST-like images. Most PE-generated images lack recognizable digits (see Gong et al. (2025)), resulting in a classification accuracy of only $\sim 30\%$ (Table 1(a)). By leveraging a simulator better suited for this domain, Sim-PE achieves significantly better results, tripling the classification accuracy and reducing the FID score by 80% at $\epsilon=10$.
Sim-PE achieves competitive results among state-of-the-art methods. When the foundation model or public data differs significantly from the private data, training-based baselines can still adapt the model to the private data distribution by updating its weights, whereas PE cannot. This limitation accounts for the substantial performance gap between PE and other methods. Specifically, PE records the lowest classification accuracy among all 12 methods (Table 1(a)). By leveraging domain-specific simulators, Sim-PE significantly narrows this gap, achieving classification accuracy within 5.4% and 4.2% of the best-performing method for $\epsilon=1$ and $\epsilon=10$, respectively.
Class label information from the simulators can be helpful. All the above experiments are based on the ClassUnavail setting, where the class label information from the simulator is assumed to be unknown. However, one key advantage of using simulators over foundation models for generating synthetic data is that simulators can provide various labels for free (Wood et al., 2021; Bae et al., 2023). In our case, for MNIST, the simulators provide information on which digit the generated image represents. Following the approach in § 4.1.1, we utilize this label information, and the results are presented in Tables 2 and 2. We observe that with digit information, the simulator-generated data achieve significantly higher classification accuracy (92.2%), although the FID remains low due to the generated digits exhibiting incorrect characteristics (Fig. 2). The fact that Sim-PE outperforms the simulator in both FID and classification accuracy across all settings suggests that Sim-PE effectively incorporates private data information to enhance both data fidelity and utility, even when compared to such a strong baseline. As expected, Sim-PE under ClassAvail matches or surpasses the results obtained in ClassUnavail across all settings, suggesting the usefulness of leveraging class label information.
(a) Real (private) images
(b) Simulator-generated images
(c) Sim-PE-generated images ($\epsilon=10$)
Table 2: Accuracy (%) of classifiers trained on synthetic images and FID of synthetic images on MNIST under the “ClassAvail” setting. See Table 1 for results under the “ClassUnavail” setting for reference.
Algorithm FID $\downarrow$ Classification Acc. $\uparrow$
$\epsilon=1$ $\epsilon=10$ $\epsilon=1$ $\epsilon=10$
Simulator 86.0 ($\epsilon=0$) 92.2 ($\epsilon=0$)
Sim-PE 20.7 8.6 93.9 95.5

### v2

@@ 3.3 Sim-PE with Simulator-generated Data
We assume a simulator-generated dataset of $m$ samples, $S_{\text{sim}}=\left\{z_{1},\ldots,z_{m}\right\}$. The goal is to select $N_{\mathrm{syn}}$ of them to form the DP synthetic dataset $S_{\mathrm{syn}}$. Before introducing our solution, we discuss why two straightforward approaches fall short.
@@
However, the size of the simulator-generated dataset (i.e., $m$) is typically very large (e.g., 1.2 million in [4]), and the total amount of added Gaussian noise grows with $m$. This means that the resulting histogram suffers from a low signal-to-noise ratio, leading to poor fidelity in $S_{\mathrm{syn}}$.
Baseline 2: Applying DP_NN_HISTOGRAM on cluster centers of $S_{\mathrm{syn}}$. To improve the signal-to-noise ratio of the histogram, one solution is to have private samples vote on the cluster centers of $S_{\text{sim}}$ instead of the raw samples. Specifically, we first cluster the samples in $S_{\text{sim}}$ into $N_{\text{cluster}}$ clusters with centers $\left\{w_{1},\ldots,w_{N_{\text{cluster}}}\right\}$ and have private samples vote on these centers rather than individual samples in $S_{\text{sim}}$. [^2] Note that voting in [46] is conducted in the image embedding space. Here, $w_{i}$s represent cluster centers in the embedding space, and each private sample uses its image embedding to find the nearest cluster center. Since the number of bins in the histogram decreases from $m$ to $N_{\text{cluster}}$, the signal-to-noise ratio improves. Following the approach of the previous baseline, we then draw $N_{\mathrm{syn}}$ cluster centers (with replacement) based on the histogram and randomly select a sample from each chosen cluster to construct the final $S_{\mathrm{syn}}$.
@@
Our approach. Our key insight is that the unavoidable trade-off between DP histogram accuracy and selection precision (clusters vs. individual samples) stems from forcing private samples to consider all of $S_{\text{sim}}$—either directly (baseline 1) or via cluster centers (baseline 2). But this is not necessary: if a sample $z_{i}$ is far from the private data, its nearest neighbors in $S_{\text{sim}}$ are likely far too (see App. C for experimental evidence). Thus, we can avoid wasting privacy budget on evaluating such samples.
@@
Since we draw only $N_{\mathrm{syn}}$ samples (instead of all $m$) from RANDOM_API, the DP histogram has a higher signal-to-noise ratio. In the following steps (Algs. 1, 1 and 1 in Alg. 1), we discard samples far from the private data and apply VARIATION_API only to the remaining ones as follows:
@@
thus avoiding consideration of nearest neighbors of the removed samples (unless they are also nearest neighbors of retained samples). Similar to $\alpha$ and $\beta$ in § 3.2, $\gamma$ controls the degree of variation. At one extreme, when $\gamma=m$, VARIATION_API disregards the input sample and reduces to RANDOM_API. At the other extreme, when $\gamma=1$, VARIATION_API returns the input sample unchanged.
Broader applications. While our main experiments (§ 4) focus on simulator-generated data, the proposed algorithm can be applied to any public dataset. We demonstrate this broader use in App. F.
@@ 3.4 Sim-PE with both Simulators and Foundation Models
As discussed in § 2.1, simulators and foundation models complement each other across different data domains. Moreover, even within a single domain, they excel in different aspects. For example, computer graphics-based face image generation frameworks [4, 61] allow controlled diversity in race, lighting, and makeup while mitigating potential biases in foundation models. However, the generated faces may appear less realistic than those produced by SoTA foundation models. Thus, combining the strengths of both methods for DP data synthesis is highly appealing.
Fortunately, PE naturally supports this integration, as RANDOM_API and VARIATION_API work the same for both foundation models and simulators. While there are many ways to combine them, we explore a simple strategy: using simulators in the early PE iterations to generate diverse seed samples, then switching to foundation models in later iterations to refine details and enhance realism. As shown in § 4, this approach outperforms using either simulators or foundation models alone.
@@ 4.1.1 Datasets and Simulators
Datasets. Following prior work [22], we use three private datasets: (1) MNIST [40], where the image class labels are digits ‘0’-‘9’, (2) CelebA [51], where the image class labels are male and female, and (3) CIFAR10 [39], which contains 10 classes of natural images. We aim at conditional generation for these datasets (i.e., each generated image is associated with the class label). Due to space constraints, CIFAR10 experiments are deferred to App. F.
Simulators. To showcase the broad applicability of Sim-PE, we use four diverse simulators.
@@
(1) Text rendering program. Generating images with readable text using foundation models is a known challenge [6]. Simulators can address this gap, as generating images with text through computer programs is straightforward. To illustrate this, we implement our own text rendering program, treating MNIST as the private dataset. Specifically, we use the Python PIL library to render digits as images. The categorical parameters include: (1) Font. We use Google Fonts [24], which offers 3589 fonts in total. (2) Text. The text consists of digits ‘0’ - ‘9’. The numerical parameters include: (1) Font size, ranging from 10 to 29. (2) Stroke width, ranging from 0 to 2. (3) Digit rotation degree, ranging from $-30\degree$ to $30\degree$. We set the feasible sets of these parameters to be large enough so that the random samples differ significantly from MNIST (see Fig. 2).
@@
(2) Computer graphics-based renderer for face images. Computer graphics-based rendering is widely used in real-world applications such as game development, cartoons, and movie production. This experiment aims to assess whether these advanced techniques can be adapted for DP synthetic image generation via Sim-PE. We use CelebA as the private dataset and a Blender-based face image renderer from [4] as the API. Since the source code for their renderer is not publicly available, we apply our data-based algorithm from § 3.3 on their released dataset of 1.2 million face images. It is important to note that this renderer may not necessarily represent the SoTA. As visualized in Fig. 3, the generated faces exhibit various unnatural artifacts and appear less realistic than images produced by SoTA generative models (e.g., [54]). Therefore, this experiment serves as a preliminary study, and the results could potentially improve with more advanced rendering techniques.
@@
(3) Rule-based avatar generator. We further investigate whether Sim-PE remains effective when the simulator’s data significantly differs from the private dataset. We use CelebA as the private dataset and a rule-based avatar generator [19] as the API. This simulator has 16 categorical parameters that control attributes of the avatar including eyes, noses, background colors, skin colors, etc. As visualized in Fig. 4, the generated avatars have a cartoon-like appearance and lack fine-grained details. This contrasts sharply with CelebA images, which consist of real human face photographs.
(4) Public images. To demonstrate that Sim-PE (§ 3.3) generalizes beyond simulator-generated data, we also evaluate it using ImageNet as $S_{\text{sim}}$. Due to space constraints, results are presented in App. F.
Class label information from simulators. In our main experiments, we assume the simulator does not provide class label information (denoted as “ClassUnavail”). In § H.3, we further investigate a setting where the simulator does provide class labels along with the generated images, and find that this leads to improved performance. See § H.3 for details and results.
@@ 4.1.2 Metrics and Evaluation Pipelines
We follow the evaluation settings of DPImageBench [22], a recent benchmark for DP image synthesis. Specifically, we use two metrics: (1) FID [31] as a quality metric and (2) the accuracy of downstream classifiers as a utility metric. Specifically, we use the conditional version of PE (App. B), so that each generated images are associated with the class labels (i.e., ‘0’-‘9’ digits in MNIST, male vs. female in CelebA). These class labels are the targets for training the classifiers. We employ a strict train-validation-test split and account for the privacy cost of classifier hyperparameter selection. Specifically, we divide the private dataset into disjoint training and validation sets. We then run Sim-PE on the training set to generate synthetic data. Next, we train three classifiers—ResNet [30], WideResNet [70], and ResNeXt [64]—on the synthetic data and evaluate their accuracy on the validation set. Since the validation set is part of the private data, we use the Report Noisy Max algorithm [16] to select the best classifier checkpoint across all epochs of all three classifiers. Finally, we report the accuracy of this classifier on the test set. This procedure ensures that the reported accuracy is not inflated due to train-test overlap or DP violations in classifier hyperparameter tuning. See App. E for further discussion on the rationale behind our choice of metrics.
Following [22], we set DP parameter $\delta=1/(N_{\mathrm{priv}}\cdot\log N_{\mathrm{priv}})$, where $N_{\mathrm{priv}}$ is the number of samples in the private dataset, and $\epsilon=1$ or 10.
@@ 4.1.3 Baselines
We compare Sim-PE with 12 SoTA DP image synthesizers reported in [22], including DP-MERF [26], DP-NTK [65], DP-Kernel [36], GS-WGAN [9], DP-GAN [63], DPDM [13], PDP-Diffusion [20], DP-LDM [50], DP-LoRA [59], PrivImage [41], and PE with foundation models [46]. Except for PE, all other baselines require model training. When experimenting with simulator-generated data, we additionally compare Sim-PE against the two baselines introduced in § 3.3.
Note: This comparison is not meant to be entirely fair, as different methods rely on different prior knowledge. For instance, many baselines use pre-trained models or public datasets from similar distributions (see Tab. 2 of [22]), whereas Sim-PE does not. Instead, Sim-PE leverages simulators, which are not used by any baseline. As such, Sim-PE represents a new evaluation setting, and baseline results serve primarily to contextualize this new paradigm and inspire future work.
@@ 4.2 Sim-PE with Simulator Access
In this section, we evaluate Sim-PE with a text rendering program on MNIST dataset. The results are shown in Tabs. 1 and 2. The key takeaway messages are:
@@
Table 1: Accuracy and FID. The best between PE methods is in bold, and the best between all methods is underlined. “Simulator” refers to samples from the simulator’s RANDOM_API. Results other than Sim-PE and Simulator are taken from [22].
@@
Quantitative results in Tab. 1 further support this. Without private data guidance, the simulator often generates digits from incorrect classes, yielding a classifier accuracy of just 11.6%–near random guess. In contrast, Sim-PE boosts accuracy to around 90%. FID scores also confirm that Sim-PE produces images more similar to real data.
Sim-PE can improve the performance of PE by a large margin. The PE baseline [46] uses a diffusion model pre-trained on ImageNet, which primarily contains natural object images (e.g., plants, animals, cars). Since MNIST differs significantly from such data, PE, as a training-free method, struggles to generate meaningful MNIST-like images. Most PE-generated images lack recognizable digits (see [22]), resulting in a classification accuracy of only $\sim 30\%$ (Tab. 1(a)). By leveraging a simulator better suited for this domain, Sim-PE achieves much better results, tripling the classification accuracy and reducing the FID by 80% at $\epsilon=10$.
Sim-PE achieves competitive results among SoTA methods. When the foundation model or public data differs significantly from the private data, training-based baselines can still adapt the model to the private data distribution by updating its weights, whereas PE cannot. This limitation accounts for the substantial performance gap between PE and other methods. Specifically, PE records the lowest classification accuracy among all 12 methods (Tab. 1(a)). By leveraging domain-specific simulators, Sim-PE substantially narrows this gap, achieving classification accuracy within 5.4% and 4.2% of the best-performing method for $\epsilon=1$ and $\epsilon=10$, respectively.